OpenStudy (anonymous):

"solve each equation, then check your solution" 5n plus 6 equals -4 Help?

OpenStudy (anonymous):

5n+6=-4 -6 -6 5n=-10 divide both sides by 5 n=-2

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

5n+6= -4 6+4 = -5n (put 5n on the other side and change the sign.) 10 = 5n (divide 10 and 5n by 5.

OpenStudy (anonymous):

Check: 5(-2)+6=-4 -10+6=-4 -4=-4